# ImageFolder: Autoregressive Image Generation with Folded Tokens

### v2

@@ 1 Introduction
VQGAN Esser et al. (2021) is a pioneer work on image tokenization that leverages a vector quantization Gray (1984) operation to map the continuous image feature to discrete tokens with a learnable codebook. Recently, several improvements from different perspectives have been made to the original VQGAN Lee et al. (2022a); Yu et al. (2023b); Mentzer et al. (2023); Zhu et al. (2024); Takida et al. (2023); Huang et al. (2023); Zheng et al. (2022); Yu et al. (2023a); Weber et al. (2024); Yu et al. (2024a); Luo et al. (2024). For example, VAR Tian et al. (2024) proposes a multi-scale vector quantization (MSVQ) which quantizes images into a series of multi-scale tokens and conducts the autoregressive modeling in a scale-based manner, leading to a remarkable latency reduction and reconstruction improvement. However, in a trade-off of the speed and reconstruction quality, the sequence length of the quantized tokens is much longer compared to other models which could result in a larger training cost and difficulty of incorporating in LLMs with long contexts Jin et al. (2024); Ding et al. (2024). SEED Ge et al. (2023) and TiTok Yu et al. (2024d) improve the VQGAN by injecting semantics into the quantized codewords. Since the semantic tokens can convey more compacted information about the image, they can achieve a much smaller token number to represent an image compared to the VQGAN. Nevertheless, a smaller token number typically leads to information loss. SEED Ge et al. (2023) tokenizer discards all pixel-level details and only retains semantics to achieve a smaller token number. TiTok Yu et al. (2024d) conducts concatenated tokenization with multiple quantization operations to achieve a higher compression rate while still suffering from detail distortion during reconstruction. Based on previous arts, we can summarize a trade-off between reconstruction and generation: (1) a longer token number can lead to a better reconstruction performance while making the sequence length too long for autoregressive generation (e.g., requiring larger model capacity Sun et al. (2024), suffering more error propagation during AR sampling Kaiser et al. (2018) and more training costs Tian et al. (2024)), and (2) a smaller token number can lead to inferior reconstruction performance while benefit generation. Motivated by this, we aim to explore whether it is possible to retain a sufficient number of tokens for high-quality reconstruction without lengthening the sequence length for generation.
@@
In this paper, we propose ImageFolder, a semantic image tokenizer that provides spatially aligned image tokens, which can be folded for a shorter token length during autoregressive modeling. We demonstrate that even if the overall token number remains the same, the folded tokens demonstrate a better generation performance compared to unfolded ones. Specifically, different from previous tokenizers Lee et al. (2022a); Yu et al. (2023b); Mentzer et al. (2023); Zhu et al. (2024); Takida et al. (2023); Huang et al. (2023); Zheng et al. (2022), we leverage product quantization to separately capture different information of images, i.e., semantics and pixel-level details, with two branches. Within each branch, we use the multi-scale residual quantization as our quantizer. A quantizer dropout is further proposed to force each residual layer to represent an image with different bitrates, thus compensating the unmodeled dependency in the next-scale AR modeling Tian et al. (2024). With the above designs, the spatially aligned tokens from ImageFolder can be folded for a parallel prediction during AR modeling as shown in Fig. 1 (b), which leads to half of the modeling token length while achieving a better generation quality. Our contribution can be summarized in three-fold:
@@ 2.1 Image Tokenizers
Image tokenizer has seen significant progress in multiple image-related tasks. Traditionally, autoencoders Hinton & Salakhutdinov (2006); Vincent et al. (2008) have been used to compress images into latent spaces for downstream work such as (1) generation and (2) understanding. In the case of generation, VAEs Van Den Oord et al. (2017); Razavi et al. (2019a) learn to map images to probabilistic distributions; VQGAN Esser et al. (2021); Razavi et al. (2019b) and its subsequent variants Lee et al. (2022a); Yu et al. (2023b); Mentzer et al. (2023); Zhu et al. (2024); Takida et al. (2023); Huang et al. (2023); Zheng et al. (2022); Yu et al. (2023a); Weber et al. (2024); Yu et al. (2024a); Luo et al. (2024) introduce a discrete latent space for better compression for generation. On the other hand, understanding tasks, such as CLIP Radford et al. (2021), DINO Oquab et al. (2023); Darcet et al. (2023), rely heavily on LLM Vaswani et al. (2023); Dosovitskiy et al. (2021) to tokenize images into semantic representations Dong et al. (2023); Ning et al. . These representations are effective for tasks like classification Dosovitskiy et al. (2021), object detection Zhu et al. (2010), segmentation Wang et al. (2021), and multi-modal application Yang et al. (2024). For a long time, image tokenizers have been divided between methods tailored for generation and those optimized for understanding. After the appearance of Yu et al. (2024c), which proved the feasibility of using LLM as a tokenizer for generation, some works Wu et al. (2024) are dedicated to unify the tokenizer for generation and understanding due to the finding in Gu et al. (2023).
@@ 3 Method
Product quantization (PQ) Jegou et al. (2010) aims to quantize a high-dimension vector to a combination of several low-dimension tokens, which has shown a promising capability to capture different contexts across sub-quantizers Baevski et al. (2020). Given a continuous feature $z$, product quantization performs as:
@@ 3.1 ImageFolder
(1) Quantizer dropout: We use multi-scale residual quantizer (MSRQ) Tian et al. (2024) for each branch, leading to multi-resolution tokens $z^{\prime}_{1},z^{\prime}_{2},\cdots,z^{\prime}_{N}$ where $N$ denotes the number of residual steps. In the $i$-th residual step, the MSRQ quantizer first downsamples the input to a smaller $L_{i}\times L_{i}$ resolution and then quantizes it by mapping it to its closest vector in the codebook $\mathcal{C}=\{e_{j}\}_{j=1}^{J}$ as $\hat{z_{i}}[a,b]=\arg\min_{e_{j}\in\mathcal{C}}\|z_{i}[a,b]-e_{j}\|_{2}^{2}$, where $[a,b]$ is the spatial coordinates. After the nearest-neighbor look-up, the quantized representation $\hat{z_{i}}$ is up-sampled back to the original resolution $K\times K$ with a convolution to handle the blurring as:
@@
where $\gamma$ is set to an constant scalar as $0.5$. Specifically, to enhance the representation capability, we notice that a quantizer dropout strategy Kumar et al. (2024); Zeghidour et al. (2021) can largely improve generation performance. During training, as shown in Fig. 5, we randomly drop out the last several quantizers. The quantizer dropout happens with a ratio of $p$. With the quantizer dropout, the final output of the MSVQ can be formulated as:
@@ 4 Experiments
3 + Smaller patch size $K=11$ 265 3.24 6.56
4 + Product quantization & Parallel decoding 265 2.06 5.96
5 + Semantic regularization on all branches 265 1.97 5.21
6 + Semantic regularization on one branch 265 1.57 3.53
7 + Strong discriminator 265 1.04 2.94
8 + Equilibrium enhancement 265 0.80 2.60
Table 3: Ablation study on the improvement of ImageFolder. We evaluate the rFID and gFID on the ImageNet validation set.
@@ 4.1 Experimental Settings
We test our ImageFolder tokenizer on the ImageNet 256x256 reconstruction and generation tasks. Our tokenizer training recipe is based on LlamaGen Sun et al. (2024). [Important] Due to the limited computational resources, we just report the results trained with a short training scheme, e.g., 250k iterations for the tokenizer. More results with advanced training recipes (more iterations, stronger discriminator, and strong perceptual loss) will be updated soon.
@@
For the ImageFolder tokenizer, if there is no other specification, we follow the VQGAN training recipe of LlamaGen Sun et al. (2024). We initialize the image encoder with the weight of DINOv2-base. We use a cosine learning rate scheduler with a warmup for 1 epoch and a start learning rate of 3e-5. We set the quantizer drop ratio to 0.1. We set $\lambda_{clip}=0.1$. We set the residual quantizer scales to $[1,1,2,3,3,4,5,6,8,11]$ (in total 265 tokens). The codebook size for each tokenizer is set to 4096. The residual quantization of each branch shares the same codebook across scales. For the generator, we adopt VAR’s Tian et al. (2024) GPT-2-based Radford et al. (2019) architecture. We double the channel of the output head to predict two tokens in parallel.
@@ 4.2 Performance analysis
We first discuss the core designs of ImageFolder. (1) We begin with our baseline using multi-scale residual quantization from VAR Tian et al. (2024) with a latent resolution of $16\times 16$ and incorporate the proposed modules step by step. (2) As shown in Tab 3, with a quantizer dropout, gFID achieves a significant improvement of 1.49 FID. We consider this to be due to the smoother next-scale AR modeling as quantizer dropout will force each scale to represent the images with different resolutions. (3) After that, we adjust the latent resolution to $11\times 1$ 11 which leads to a performance drop for both reconstruction and generation due to a smaller token number. (4) To increase the token number without enlarging the token length, we adopt a product quantization with corresponding parallel decoding. This modification results in a rFID of 2.06 and a gFID of 5.96 which outperforms the $16\times 16$ counterpart’s gFID of 6.03. (5) We further employ semantic regularization to enrich the semantics in the latent space for a more compact representation. (6) With the analysis of Tab 2, we realize that pure semantics is not enough for image reconstruction. Therefore, we adjust the semantic regularization to only one branch of the product quantization. Significant performance improvement is observed in both reconstruction and generation. Regularization on a single branch encourages different branches to capture different information about the image for a better reconstruction quality and reduces the feature dependency thus leading to a more robust sampling during generation. With all components, the VAR with ImageFolder achieves 3.53 gFID with only 265 token lengths.
@@
Diff. LDM-4 Rombach et al. (2022b) 0.74 - 3.60 247.7 - - 400M - 250
Diff. DiT-L/2 Peebles & Xie (2023b) - - 5.02 167.2 0.75 0.57 458M - 250
@@
AR. ImageFolder (ours) 0.80 58.0 2.60 295.0 0.75 0.63 362M 265 10
Table 4: Performance comparison on class-conditional ImageNet 256x256. [Important] Due to the computing limitation, the utilized ImageFolder tokenizer is trained with a short scheme, i.e., 250K iterations. Results with advanced training recipes as other comparable methods will be updated soon. denotes the utilized image tokenizer is trained on OpenImage Kuznetsova et al. (2020).
@@
As shown in Tab 4, we compare ImageFolder with state-of-the-art generative models. We notice ImageFolder outperforms LlamaGen which shares a similar tokenizer training scheme. Compared with VAR (with advanced tokenizer training recipe, e.g., OpenImage, longer training, and stronger discriminator), our method achieves a comparable performance. In addition, the shorter token length of ImageFolder (265 length) also remarkably saves training/inference costs compared to VAR (680 length) which is $\mathcal{O}(n^{2})$ to the length $n$. Since VAR Tian et al. (2024) did not release their tokenizer training code and recipe, we can not reproduce their result. With our architecture and training recipe, the vanilla MSVQ Tian et al. (2024) only leads to a rFID of 1.92 and a gFID of 7.52.
@@ 4.3 More Analysis
Method LlamaGen VAR ImageFolder
Acc. 2.62 11.32 55.42
Table 5: Linear probing Top-1 accuracy comparison on ImageNet val set.

### v3

@@ 1 Introduction
VQGAN Esser et al. (2021) is a pioneer work on image tokenization that leverages a vector quantization Gray (1984) operation to map the continuous image feature to discrete tokens with a learnable codebook. Recently, several improvements from different perspectives have been made to the original VQGAN Lee et al. (2022a); Yu et al. (2023b); Mentzer et al. (2023); Zhu et al. (2024a); Takida et al. (2023); Huang et al. (2023); Zheng et al. (2022); Yu et al. (2023a); Weber et al. (2024); Yu et al. (2024a); Luo et al. (2024). For example, VAR Tian et al. (2024) proposes a multi-scale vector quantization (MSVQ) which quantizes images into a series of multi-scale tokens and conducts the autoregressive modeling in a scale-based manner, leading to a remarkable latency reduction and reconstruction improvement. However, in a trade-off of the speed and reconstruction quality, the sequence length of the quantized tokens is much longer compared to other models which could result in a larger training cost and difficulty of incorporating in LLMs with long contexts Jin et al. (2024); Ding et al. (2024). SEED Ge et al. (2023) and TiTok Yu et al. (2024d) improve the VQGAN by injecting semantics into the quantized codewords. Since the semantic tokens can convey more compacted information about the image, they can achieve a much smaller token number to represent an image compared to the VQGAN. Nevertheless, a smaller token number typically leads to information loss. SEED Ge et al. (2023) tokenizer discards all pixel-level details and only retains semantics to achieve a smaller token number. TiTok Yu et al. (2024d) conducts concatenated tokenization with multiple quantization operations to achieve a higher compression rate while still suffering from detail distortion during reconstruction. Based on previous arts, we can summarize a trade-off between reconstruction and generation: (1) a longer token number can lead to a better reconstruction performance while making the sequence length too long for autoregressive generation (e.g., requiring larger model capacity Sun et al. (2024), suffering more error propagation during AR sampling Kaiser et al. (2018) and more training costs Tian et al. (2024)), and (2) a smaller token number can lead to inferior reconstruction performance while benefit generation. Motivated by this, we aim to explore whether it is possible to retain a sufficient number of tokens for high-quality reconstruction without lengthening the sequence length for generation.
@@
In this paper, we propose ImageFolder, a semantic image tokenizer that provides spatially aligned image tokens, which can be folded for a shorter token length during autoregressive modeling. We demonstrate that even if the overall token number remains the same, the folded tokens demonstrate a better generation performance compared to unfolded ones. Specifically, different from previous tokenizers Lee et al. (2022a); Yu et al. (2023b); Mentzer et al. (2023); Zhu et al. (2024a); Takida et al. (2023); Huang et al. (2023); Zheng et al. (2022), we leverage product quantization to separately capture different information of images, i.e., semantics and pixel-level details, with two branches. Within each branch, we use the multi-scale residual quantization as our quantizer. A quantizer dropout is further proposed to force each residual layer to represent an image with different bitrates, thus compensating the unmodeled dependency in the next-scale AR modeling Tian et al. (2024). With the above designs, the spatially aligned tokens from ImageFolder can be folded for a parallel prediction during AR modeling as shown in Fig. 1 (b), which leads to half of the modeling token length while achieving a better generation quality. Our contribution can be summarized in three-fold:
@@ 2.1 Image Tokenizers
Image tokenizer has seen significant progress in multiple image-related tasks. Traditionally, autoencoders Hinton & Salakhutdinov (2006); Vincent et al. (2008) have been used to compress images into latent spaces for downstream work such as (1) generation and (2) understanding. In the case of generation, VAEs Van Den Oord et al. (2017); Razavi et al. (2019a) learn to map images to probabilistic distributions; VQGAN Esser et al. (2021); Razavi et al. (2019b) and its subsequent variants Lee et al. (2022a); Yu et al. (2023b); Mentzer et al. (2023); Zhu et al. (2024a); Takida et al. (2023); Huang et al. (2023); Zheng et al. (2022); Yu et al. (2023a); Weber et al. (2024); Yu et al. (2024a); Luo et al. (2024); Zhu et al. (2024b) introduce a discrete latent space for better compression for generation. On the other hand, understanding tasks, such as CLIP Radford et al. (2021), DINO Oquab et al. (2023); Darcet et al. (2023); Zhu et al. (2024c), rely heavily on LLM Vaswani et al. (2023); Dosovitskiy et al. (2021) to tokenize images into semantic representations Dong et al. (2023); Ning et al. . These representations are effective for tasks like classification Dosovitskiy et al. (2021), object detection Zhu et al. (2010), segmentation Wang et al. (2021), and multi-modal application Yang et al. (2024). For a long time, image tokenizers have been divided between methods tailored for generation and those optimized for understanding. After the appearance of Yu et al. (2024c), which proved the feasibility of using LLM as a tokenizer for generation, some works Wu et al. (2024) are dedicated to unify the tokenizer for generation and understanding due to the finding in Gu et al. (2023).
@@ 3 Method
Product quantization (PQ) Jegou et al. (2010); Guo et al. (2024) aims to quantize a high-dimension vector to a combination of several low-dimension tokens, which has shown a promising capability to capture different contexts across sub-quantizers Baevski et al. (2020). Given a continuous feature $z$, product quantization performs as:
@@ 3.1 ImageFolder
(1) Quantizer dropout: We use multi-scale residual quantizer (MSRQ) Tian et al. (2024) for each branch, leading to multi-resolution tokens $z^{\prime}_{1},z^{\prime}_{2},\cdots,z^{\prime}_{N}$ where $N$ denotes the number of residual steps. In the $i$-th residual step, the MSRQ quantizer first downsamples the input to a smaller $K_{i}\times K_{i}$ resolution and then quantizes it by mapping it to its closest vector in the codebook $\mathcal{C}=\{e_{j}\}_{j=1}^{J}$ as $\hat{z_{i}}[a,b]=\arg\min_{e_{j}\in\mathcal{C}}\|z_{i}[a,b]-e_{j}\|_{2}^{2}$, where $[a,b]$ is the spatial coordinates. After the nearest-neighbor look-up, the quantized representation $\hat{z_{i}}$ is up-sampled back to the original resolution $K\times K$ with a convolution to handle the blurring as:
@@
where $\gamma$ is set to a constant scalar as $0.5$. Specifically, to enhance the representation capability, we notice that a quantizer dropout strategy Kumar et al. (2024); Zeghidour et al. (2021) can largely improve generation performance. During training, as shown in Fig. 5, we randomly drop out the last several quantizers. The quantizer dropout happens with a ratio of $p$. With the quantizer dropout, the final output of the MSVQ can be formulated as:
@@ 4 Experiments
3 + Smaller quantized token size $K=11$ 286 3.24 6.56
4 + Product quantization & Parallel decoding 286 2.06 5.96
5 + Semantic regularization on all branches 286 1.97 5.21
6 + Semantic regularization on one branch 286 1.57 3.53
7 + DINO discriminator 286 0.80 2.60
Table 3: Ablation study on the improvement of ImageFolder. We evaluate the rFID and gFID on the ImageNet validation set. DINO discriminator denotes the discriminator used in VAR’s Tian et al. (2024) tokenizer. $\uparrow$ and $\downarrow$ denote the larger the better and the smaller the better, respectively.
@@ 4.1 Experimental Settings
We test our ImageFolder tokenizer on the ImageNet 256x256 reconstruction and generation tasks. Our tokenizer training recipe is based on LlamaGen Sun et al. (2024). More results with advanced training recipes (more iterations, stronger discriminator, and strong perceptual loss) will be updated during the author-reviewer discussion.
@@
For the ImageFolder tokenizer, if there is no other specification, we follow the VQGAN training recipe of LlamaGen Sun et al. (2024). We initialize the image encoder with the weight of DINOv2-base. We use a cosine learning rate scheduler with a warmup for 1 epoch and a start learning rate of 3e-5. We set the quantizer drop ratio to 0.1. We set $\lambda_{clip}=0.1$, $\lambda_{recon}=\lambda_{VQ}=\lambda_{P}=1$ and $\lambda_{ad}=0.5$. We set the residual quantizer scales to $[1,1,2,3,3,4,5,6,8,11]$ (in total 286 tokens). The codebook size for each tokenizer is set to 4096. The residual quantization of each branch shares the same codebook across scales. For the generator, we adopt VAR’s Tian et al. (2024) GPT-2-based Radford et al. (2019) architecture. We double the channel of the output head to predict two tokens in parallel.
@@ 4.2 Performance analysis
We first discuss the core designs of ImageFolder. (1) We begin with our baseline using multi-scale residual quantization from VAR Tian et al. (2024) with a latent resolution of $16\times 16$ and incorporate the proposed modules step by step. (2) As shown in Tab 3, with a quantizer dropout, gFID achieves a significant improvement of 1.49 FID. We consider this to be due to the smoother next-scale AR modeling as quantizer dropout will force each scale to represent the images with different resolutions. (3) After that, we adjust the latent resolution to $11\times 1$ 11 which leads to a performance drop for both reconstruction and generation due to a smaller token number. (4) To reduce the token length without decreasing the token number, we adopt a product quantization with corresponding parallel decoding. This modification results in a rFID of 2.06 and a gFID of 5.96 which outperforms the $16\times 16$ counterpart’s gFID of 6.03. (5) We further employ semantic regularization to enrich the semantics in the latent space for a more compact representation. (6) With the analysis of Tab 2, we realize that pure semantics is not enough for image reconstruction. Therefore, we adjust the semantic regularization to only one branch of the product quantization. Significant performance improvement is observed in both reconstruction and generation. Regularization on a single branch encourages different branches to capture different information about the image for a better reconstruction quality and reduces the feature dependency thus leading to a more robust sampling during generation. With all components, the VAR with ImageFolder achieves 3.53 gFID with only 286 token lengths.
@@
Diff. LDM-4 Rombach et al. (2022b) - - 3.60 247.7 - - 400M - 250
Diff. DiT-L/2 Peebles & Xie (2023b) 0.90 - 5.02 167.2 0.75 0.57 458M - 250
@@
AR. ImageFolder (ours) 0.80 58.0 2.60 295.0 0.75 0.63 362M 286 10
Table 4: Performance comparison on class-conditional ImageNet 256x256. Results with advanced training recipes as other comparable methods will be updated soon. denotes the utilized image tokenizer is trained on OpenImage Kuznetsova et al. (2020). $\uparrow$ and $\downarrow$ denote the larger the better and the smaller the better, respectively. L.P. denotes linear probing accuracy on the ImageNet val set. The codebook utilization rate for the Imagefolder tokenizer is 100%.
@@
As shown in Tab 4, we compare ImageFolder with state-of-the-art generative models. We notice ImageFolder outperforms LlamaGen which shares a similar tokenizer training scheme. Compared with VAR (with advanced tokenizer training recipe, e.g., OpenImage, longer training, and stronger discriminator), our method achieves a comparable performance. In addition, the shorter token length of ImageFolder (286 length) also remarkably saves training/inference costs compared to VAR (680 length) which is $\mathcal{O}(n^{2})$ to the length $n$. Since VAR Tian et al. (2024) did not release their tokenizer training code and recipe, we can not reproduce their result. With our architecture and training recipe, the vanilla MSVQ Tian et al. (2024) only leads to a rFID of 1.92 and a gFID of 7.52.
@@ Appendix D Teacher Forcing Guidance for Conditional Generation
Appendix E Additional Experiments
ID Branch1 Branch2 rFID gFID
1 - - 2.06 5.96
2 DINOv2 DINOv2 1.97 5.96
3 DINOv2 - 1.57 3.53
4 DINOv2 SAM 1.47 4.05
Table 7: Ablation study on regularization.
Regularization on more branches.
We demonstrate the ablation study to investigate the performance of different encoders to conduct distillation. We notice that leveraging different encoders, i.e., DINO & SAM, to conduct regularization can lead to a superior rFID. However, we notice that the gFID does not improve as the rFID. Instead, it shows an inferior performance.
Length Scales rFID gFID
265 1,1,2,3,3,4,5,6,8,10 1.86 3.98
286 1,1,2,3,3,4,5,6,8,11 1.57 3.53
309 1,1,2,3,3,4,5,6,8,12 1.22 3.51
Table 8: Ablation study on token length.
Token length.
We demonstrate an additional ablation study on token length where the patch size of the last scale increases from 10 to 12 leading to 265 to 309 token length for autoregressive modeling. We notice that the rFID keeps decreasing when the token length increases while the gFID is saturated for a token length near 286. Therefore, we utilize 286 token lengths in our final setting.
Codebook size 4096 8192 16384
rFID 0.80 0.70 0.67
Table 9: Ablation study on codebook size.
Token length.
We conduct an ablation study on the codebook size of the ImageFolder tokenizer. We noticed that a large codebook size of 16384 makes a great improvement to the rFID to 0.67 which is comparable to the continuous tokenizers.
Method Branch rFID gFID
Single detail branch 1 3.24 6.56
Single semantic branch 1 2.79 5.81
Two semantic branches 2 1.97 5.21
Two detail branches 2 2.06 5.96
One semantic, one detail 2 1.57 3.53
Table 10: Comparison of different methods with the corresponding branch, rFID, and gFID values.
Branch setting.
We provide additional experiments complimentary to Table 3 to analyze the branch settings. We notice that in the setting of two branches with one for semantic information and one for detail information, the model achieves the best performance.
Figure 13: A visualization of a product quantized space with 2 subspaces.
Appendix F Product Quantized Space
Product quantization divides the original space into subspaces and then conducts quantization on the subspaces. As shown in Fig. 13, we demonstrate two scenarios of product quantized space with different data sample layouts.
As the product quantization leverages the cartesian product to connect the separately quantized subspaces, it will ensure a symmetric property of the quantized space (symmetric indicates the quantized tokens are symmetric to the axes, showing a grid pattern). In this way, if the original data sample is strictly asymmetric, considering a two-subspace case (Fig. 13), it will require 2$\times$##\##sample codewords to perfectly encode all the samples. Differently, if the data samples are arranged in a symmetric manner, the required number of codewords can be largely reduced to $2\times\sqrt{\text{$\#$sample}}$.
In this way, a potential improvement direction is to find subspaces that have desired properties to be quantized. In the ImageFolder framework, this can be achieved by leveraging contrastive loss to guide the latent space in different branches.